OpenStudy (anonymous):

. The price of a ticket for a certain concert is $5 for a child, $7 for a student, and $10 for an adult. Suppose that there were 260 people in attendance and that the total amount received in ticket sales was $1,932. Moreover, suppose that there were three times as many children at the concert as there were adults. How many children, students, and adults were at the concert? Find a system of linear equations which models this problem. Be sure to clearly identify the variables used. DO NOT SOLVE THE SYSTE

OpenStudy (anonymous):

Na = number of adults. Nc = number of child Ns = number of students Here are the 3 equations : \[Nc = 3 Na\] 260 = Nc + Na + Ns \[260 = 4 Na + Ns\] 1932 = 5 Nc + 7 Ns + 10 Na 1932 = 5 * 3 Na + 7 Ns + 10 Na \[1932 = 7 Ns + 25 Na\]